OpenStudy (anonymous):

The initial start-up cost for computing equipment, facilities, course development, and staff recruitment and development is $350,000. The college plans to charge tuition of $18,000 per student per year. However, the university administration will charge the college $12,000 per student for the first 100 students enrolled each year for administrative costs and its share of the tuition payments. a. How many students does the college need to enroll in the first year to break even? b. If the college can enroll 75 students the first year, how much profit will it make?

OpenStudy (anonymous):

please help :((

OpenStudy (anonymous):

11 1900000 in profit

OpenStudy (anonymous):

hope it helps

OpenStudy (anonymous):

can you give me some clue on how to get the cost function? thanks you so much:)

OpenStudy (kropot72):

Let the number of students needed to break even = S Annual income = 18000 * S Annual expenses = 350000 + (12000 * S) To break even the annual income must equal the annual expenses: 18000S = 350000 + 12000S 6000S = 350000 Now solve for S to find the number of students to break even. Can you do that?

OpenStudy (anonymous):

yes, i know the annual income, i am not that sure on the fixed and constant cost for annual expenses, thanks for that:))

OpenStudy (kropot72):

Obviously the initial start-up cost must be recovered. The question does not state a recovery period, therefore it can be assumed that all the start-up cost will be recovered in the first year.

OpenStudy (anonymous):

i see, thanks again :)

OpenStudy (kropot72):

You're welcome :)